OpenStudy (rahulmr):

given \[f(x)=k*(k-1)^2*e^{-x}+k\] a) For what values of x in terms of k will f(x) have exactly 2 stationary points. b) Find the exact co-ordinates of local max point. Ok. So I believe we differentiate f(x) and make f'(x)=0 and solve for k. I got k=0 and k=1. But I don't know if its right or not. Can anyone help me out. Thanks.

OpenStudy (3mar):

I can help with graph if you did not try that way.

OpenStudy (3mar):

Feel free to try www.calorful.com and manipulate your constant k as you wish.

OpenStudy (rahulmr):

Thanks for the reply, but could you check my way is correct though, because in the next question, it asks to find the local max and if the graph is a straight line, there is no max.

OpenStudy (3mar):

Correct! No max for a straight line. Let me check your way.

OpenStudy (3mar):

So what are your steps? could you share them with me?

OpenStudy (rahulmr):

I differentiate f(x) and then make it equal to 0 to solve for k. So I got k=0 and k=1. But i think you sub it back into the equation, however it does not makes sense.

zepdrix (zepdrix):

They're asking for values of x though, not values of k. Hmm this is weird :p

OpenStudy (3mar):

I have done what you said, but I got nothing!

zepdrix (zepdrix):

Like.. umm.. if x=ln(k), then,\[\large\rm f'=-k(k-1)^2e^{-\ln k}\]\[\large\rm f'=-k(k-1)^2e^{\ln \frac1k}\]\[\large\rm f'=-k(k-1)^2\frac1k\]\[\large\rm f'=(k-1)^2\]

zepdrix (zepdrix):

Honestly I don't know if that's what they're looking for though... this problem doesn't make much sense to me :(

zepdrix (zepdrix):

Woops that last line should be,\[\large\rm f'=-(k-1)^2\]

OpenStudy (3mar):

why \[x=\ln(k)\]?

OpenStudy (rahulmr):

i have\[-k*(k-1)^2*e^{-x}\]

OpenStudy (3mar):

yes, e^(-x) not e^(-ln(k))